Jack went to the cash point. He drew out the same amount of money as he had in his pocket. He then spent £40 on CDs.
Jack went back to the cash point.
He drew out the same amount of money as he had in his pocket.
He then spent a further £40 on CDs. Jack did this a third time.
He then had no money left. How much money did he have with him to start?
1
Expert's answer
2020-10-14T18:48:31-0400

Suppose that Jack had with him to start £x.

Jack went to the cash point. He drew out the same amount of money as he had in his pocket.


"\u00a3x+\u00a3x=\u00a3(2x)"

He then spent £40 on CDs


"\u00a3(2x-40)"

Jack went back to the cash point. He drew out the same amount of money as he had in his pocket.


"2\u00a3(2x-40)=\u00a3(4x-80)"

He then spent a further £40 on CDs


"\u00a3(4x-80-40)=\u00a3(4x-120)"

Jack went back to the cash point. He drew out the same amount of money as he had in his pocket.


"2\u00a3(4x-120)=\u00a3(8x-240)"

He then spent a further £40 on CDs


"\u00a3(8x-240-40)=\u00a3(8x-280)"

He then had no money left


"\u00a3(8x-280)=\u00a3(0)"

"8x-280=0"

"8x=280"

"x=35"

Jack had with him to start £35.